OpenStudy (anonymous):

The Office Jet printer can copy Lisa's dissertation in 12 min. The Laser Jet printer can copy the same document in 20 min. If the two machines work together, how long would they take to copy the dissertation?

OpenStudy (anonymous):

is the answer 8?

OpenStudy (anonymous):

The combined rate is: 1/20 + 1/12 = (6 + 10)/120 = 1/7.5 dissertations per minute. (1/7.5)*T = 1 dissertation T = 7.5 minutes

OpenStudy (anonymous):

thank you!